# ConceptDrift: Uncovering Biases through the Lens of Foundation Models

### v2

@@ 4 Experimental Setup
Foundation models (FM) We used mGTE (gte-large-en-v1.5 [45]) for text embeddings in CivilComments [5], and OpenAI CLIP ViT-L/14 [30] for text and images otherwise.
Concept extraction For the image classification task, we first use the GIT-Large  [41] captioning model (trained on MSCOCO [19]) to obtain descriptions of the dataset’s images. Next, we extract concepts from the captions (or directly from the text samples for the text classification task), using YAKE [7] keyword extractor, taking the top $256$ n-grams for $n=3,5$.
Concepts filtering We use Llama-3.1-8B-Instruct [11] to remove class instances from the concepts extracted at the previous step. We also apply post-processing using WordNet [24] to catch obvious class instances that the LLM might miss. To do this, for each class we specify a name used to search for synsets in the WordNet [24] hierarchy (e.g. bird for Waterbirds) and then remove individual words that match with any hyponym or hypernym of those synsets.
Training We train the linear layer on $L_{2}$ normalized embeddings extracted by these models using the PyTorch [26] AdamW [23] optimizer with a learning rate of $41e-$ 4, a weight decay of $51e-$ 5 and a batch size of $1024$. We use the cross entropy loss with balanced class weights as the objective. The weights of the layer are normalized after each update and we use CLIP’s [30] temperature to scale the logits. We use the class-balanced accuracy on the validation set as the criterion for model selection and early stopping.
@@ 4.1 Datasets
Waterbirds [40] is a common datasets for generalization and bias mitigation. It is created from CUB [39], by grouping species of birds into two categories, landbirds and waterbirds, each being associated with a spurious correlation regarding its background, land and water respectively.
@@
CelebA [21] is a large-scale collection of celebrity images (over 200000), widely used in computer vision research. The setup for using it in a generalization context [22] consists of using the Blond_Hair attribute as the class label and the gender as the spurious feature.
@@
CivilComments [5] is a large collection comprised of $1.8$ million online user comments. This dataset is used employed in NLP bias and fairness research concerning different social and ethnical groups.
@@
ImageNet [10] is a larger-scale popular dataset for image classification ($1000$ classes, with approximately $1300$ training samples and $50$ validation samples per class).
5 Bias identification evaluation
@@
For a proper quantitative evaluation of our proposed biases, beyond the subjectivity of the qualitative aspects, we combine the bias identified by ConceptDrift with different components: in Sec. 5.1 we generate samples exploiting the discovered biases; in Sec. 5.2 we assess GroupDRO [33] robustness over training set partitions based on the proposed biases groups; in Sec. 5.3 we evaluate bias-based training regularization; and in Sec. 5.4 we use a minimally bias-enhanced prompting for zero-shot prediction with a Foundational Model. An exhaustive list of the extracted biases for all datasets can be found in Appx. B.
5.1 ImageNet biases
We employ ConceptDrift to uncover a series of biases and spurious correlations plaguing the decision-making process of OpenAI’s CLIP ViT-L/14 fine-tuned on ImageNet. As the biases our method reveals lie beyond the restricted scope of the validation set, our sole recourse for illustrating and validating these findings at scale lies in generating synthetic images. To this end, we have employed a quantized version of FLUX.1-dev [17]. To validate the impact of the biases, we prompt the generative model to depict a chosen (correct) class alongside a non-ImageNet object that we identified as a positively correlated bias for a different class.
The validation process is presented for three distinct scenarios in Tab. 2. Each scenario is defined by a pair of classes: a correct class which is illustrated in the image and another class which is absent, but induced through the presence of a biasing object or activity. We measure the impact of the bias by means of comparing the model’s ability to predict the correct class before and after the introduction of the biasing factor. We generate $1000$ images for each prompt, and we evaluate the model both in terms of the predicted class and in terms of the activations of the class neurons. We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved, and a significantly increased likelihood to have the induced class predicted, even though it is not illustrated in any way, shape or form in the generated image.
Within the same context, we present a series of qualitative examples in Fig. 3. We emphasize that, even though throughout most of these samples, a single ImageNet class is clearly depicted, the model chooses to ignore it and label the image as a completely different class, not illustrated at all in the image, solely based on the presence of a non-ImageNet object. As an example, the network chooses to completely ignore the peafowl, present right at the center of the second image, and labels the image as a ”Fire Truck” based on the presence of the firefighters.
The model at hand is generally considered to be a robust state-of-the-art model, benefiting from ample pre-training, achieving an $82.4\%$ top-1 accuracy on the ImageNet validation set. We emphasize through our experiment that even under these reassuring circumstances, critical reasoning flaws can make their way through in a production-ready model, undetected by validating and examining the model’s performance on held-out data (the traditional approach for identifying biases). It is through shifting the approach towards investigating the model’s weights that we can anticipate and prevent serious biases from making their way into the final version of the model.
Throughout this validation process, we have manually ensured that the generated data does not contain troubling artifacts. For instance, it is the case that even though the prompt only asks for a firefighter, in certain circumstances the generative network will also depict a fire truck alongside the other features requested by the prompt, against our desires. Additionally, sometimes the generative model is not precise and may depict a dog or a bird species different than the one we asked for. These examples of faulty generation are not exhaustive. It is therefore the case that prompts have to be carefully designed and images carefully inspected, in order to ensure the quality of the evaluation. Under these circumstances, we have endeavored to illustrate and validate but a small set of the remarkable flaws, ascribing a complete examination of our findings to future work. However, the complete list of ranked concepts for each class, determined automatically by employing our ConceptDrift approach, is available for consultation in Appx. B.
@@ 5.1 ImageNet biases
Table 3: Learning in the context of 100% spurious correlations. In the absence of samples with negatively correlated biases, GroupDRO performs similarly with the standard ERM. In contrast, bias regularization using our biases yields excellent outcomes in this setup, improving the worst group accuracy above the considered baselines: ERM, GroupDRO and the bias regularization with random causally-unrelated concepts (obtained after filtering in Step2a.).
@@
Bias regularization w/ random concepts (class-neutral) 46.6$\pm$2.7 75.3$\pm$1.1 9.4$\pm$0.0 61.4$\pm$2.0 19.1$\pm$1.6 50.8$\pm$0.9
Bias regularization w/ ConceptDrift biases (ours) 57.9$\pm$0.3 79.8$\pm$0.1 10.4$\pm$0.5 62.0$\pm$1.8 31.3$\pm$0.7 57.5$\pm$0.4
5.2 From biases to groups, towards generalization
Next, we move to the subpopulation shift setup [33] (where different biases determine subgroups in the data) and follow the evaluation protocol in SubpopBench [43]. Normally, ERM training on such data behaves poorly on minority groups, but methods like GroupDRO [33] can recover the performance, when the subgroups are given. We show that the biases discovered by our method can be used to create such subgroups, proving that they are indeed spuriously correlated concepts that a robust model should not rely on.
Table 4: Generalization task, evaluated by comparing the performance of an ImageNet-pretrained ResNet-50 model, trained with GroupDRO on dataset partitions that are based on the identified biases. ConceptDrift biases are competitive with SoTA.
. Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) Training Bias identification Worst Avg. Worst Avg. ERM - 70.0$\pm$2.4 84.9$\pm$0.9 54.2$\pm$3.6 95.4$\pm$0.3 GroupDRO w/ B2T [15] partitioning into groups B2T [15] 83.2±$\pm 3.6$ 90.6±$\pm 2.0$ 87.6±$\pm 2.0$ 93.7±$\pm 0.4$ Lg [47] 81.4±$\pm 3.4$ 89.5±$\pm 1.3$ 89.3±$\pm 1.8$ 93.2±$\pm 0.6$ ConceptDrift (ours) 81.9±$\pm 2.2$ 90.0±$\pm 2.3$ 88.4±$\pm 1.4$ 93.6±$\pm 0.4$ Ground-truth groups 85.8±$\pm 2.5$ 90.1±$\pm 1.6$ 88.8±$\pm 2.0$ 93.4±$\pm 0.7$
For the experiments in Tab. 4, we use the following pipeline. Based on the identified biases, we split the training data into subgroups with the partitioning technique from B2T [15]. Next, we train the model with GroupDRO, constraining it to attend equitably to the subgroups. If the concepts represent biases that will lead the model to learn spurious correlations through ERM, then under GroupDRO training, the model will reduce its reliance on them, surpassing the performance of ERM on the minority subgroups.
Tab. 4 shows that our identified ConceptDrift-biases are indeed useful for GroupDRO, far surpassing the ERM baseline and obtaining results that are competitive with the other bias identification methods in terms of worst-group performance on the Waterbirds and CelebA datasets. Note that the partitioning method was not tuned to accommodate our biases (e.g. our biases include actions and attributes, while the other methods only propose nouns).
Loss correlates with the presence of biases In the setup described above, we look at the correlations between the loss values and the bias-to-sample similarities. We compare ERM with GroupDRO applied on subgroups obtained with our biases and the B2T partitioning strategy. See in Appx. D how after GroupDRO with our proposed biases, the correlations significantly decrease, revealing that the model is less prone to make mistakes on the biased samples. Intuitively, this also indicates that the model is less reliant on the biases.
5.3 Biases used for training regularization
We previously showed in Sec. 5.2 that partitioning datasets using qualitative biases, followed by GroupDRO training, can significantly improve the robustness of a model to the bias-induced distribution shifts. We now investigate a scenario where prior methods for robustness like GropuDRO or others [33, 44, 14] are not applicable. In a setup with 100% bias-to-class correlations, the absence of the diverse bias-class associations renders these existing methods ineffective. We thus propose the task of learning in this perfect correlated setup, removing all bias conflicting samples from the training and validation sets.
We use the discovered biases to impose regularization upon the trained models, in particular, regularization for linear probing, learned on top of frozen CLIP embeddings. Intuitively, we constraint each bias to be equally distanced from the class weights. We formulate this as a MSE between the similarity of $b$ with a class weight and the average similarity of $b$ to all the class weights, scaled by $\tau$ (CLIP’s temperature):
@@ 5.3 Biases used for training regularization
where $\mathcal{B}$ is the set of selected biases and $\alpha=0.1$.
In Tab. 3, we present the results of linear probing with $\mathcal{L}$ loss, in this perfect spurious correlation scenario. The improvement in worst group accuracy shows that the learned classification weights are less reliant on the biases, thus more robust and better aligned with the classification task.
Table 5: Results for the Foundation Model (FM) Zero-shot prompting task. We help the model account for biases by explicitly introducing them in the prompting scheme. We perform tests employing biases discovered by different methods and evaluate the zero-shot performance of the FM. Aided by the biases discovered through our proposed method, the model’s performance, measured in terms of worst group accuracy, is significantly enhanced across all test scenarios. Our method surpasses in this regard all previous bias-identification approaches.
@@
FM zero-shot 35.2 90.7 72.8 87.4 33.1 83.4
FM w B2T [15] 48.1 86.1 72.8 88.0 - -
FM w SpLiCE [4] - - 67.2 90.2 - -
FM w Lg [47] 46.1 85.9 50.6 87.2 - -
FM w ConceptDrift 50.3 86.3 73.1 85.7 53.2 71.0
Table 6: Ablation. Following the zero-shot bias-augmented prompting setup, we variate the cut-off threshold for considering causally-unrelated concepts as biases and also try to address the text-image modality gap.
@@
5.4 Bias-aware zero-shot prompting
To further validate our identified biases, we follow Kim et al. [15] and evaluate them in the context of zero-shot classification task. We augment the initial, class-only, related prompt with the bias only through a minimal intervention (e.g.’a photo of a {cls} in the {bias}’ (see Appx. A). For each class we create a prompt with each identified bias. When classifying an image, we take into account only the highest similarity among the prompts of a class (zero-shot with max-pooling over prompts). The results in Tab. 5 show how the biases automatically selected by our method improve the worst group accuracy over the initial zero-shot baseline and other state-of-the-art bias-extracting solutions in all the tested datasets. This highlights the relevance of the biases automatically extracted by ConceptDrift.
Ablations We validate several key decisions in our algorithm in Tab. 6, using the zero-shot classification task and the biased-enhanced prompting. We notice that the number of chosen class-unrelated concepts is important, as taking too many of them adds noise to the prompts and lowers the performance. Nevertheless, dynamically choosing the threshold, as described in Sec. 3-Step2c., proves to be a good strategy for adapting the cut-off across datasets. Following prior observations regarding the modalities gap between text and image embedding space [18, 46], we follow Liang et al. [18] and subtract half of the gap from the embeddings, then re-normalize them, ending in marginally lower performance w.r.t. not addressing the modality gap.
@@ 6 Conclusions
We introduce ConceptDrift, a method that identifies biases through a weight-space and filtration approach, that can be scaled to large datasets. By leveraging CLIP models with aligned image-text embedding, we can easily identify concepts that significantly influence class predictions, while filtering undesirable concepts. We apply this method to multiple modalities on datasets like Waterbirds, CelebA, CivilComments, and ImageNet. We validate that the desired biases are useful by employing them in different methods for improved generalization. This includes a) creating subgroups that promote robust generalization b) a novel bias regularization that works even under $100\%$ correlation of biases and classes, where previous methods fail c) using better bias-aware zero-shot promoting. Using ConceptDrift we identify biases in ImageNet and generate synthetic images that are able to fool strong pre-trained methods, validating the relevance of the identified biases in larger datasets. This method opens the door for automatically finding biases in large settings without explicit human intervention.
Supplementary Material

### v3

@@ 4 Experimental Setup
Foundation models (FM) We used mGTE (gte-large-en-v1.5 (Zhang et al., 2024)) for text embeddings in CivilComments (Borkan et al., 2019), and OpenAI CLIP ViT-L/14 (Radford et al., 2021) for text and images otherwise.
Concept extraction For the image classification task, we first use the GIT-Large  (Wang et al., 2022) captioning model (trained on MSCOCO (Lin et al., 2014)) to obtain descriptions of the dataset’s images. Next, we extract concepts from the captions (or directly from the text samples for the text classification task), using YAKE (Campos et al., 2020) keyword extractor, taking the top $256$ n-grams for $n=3,5$.
Concepts filtering We use Llama-3.1-8B-Instruct (Dubey et al., 2024) to remove class instances from the concepts extracted at the previous step. We also apply post-processing using WordNet (Miller, 1995) to catch obvious class instances that the LLM might miss. To do this, for each class we specify a name used to search for synsets in the WordNet (Miller, 1995) hierarchy (e.gbird for Waterbirds) and then remove individual words that match with any hyponym or hypernym of those synsets.
Training We train the linear layer on $L_{2}$ normalized embeddings extracted by these models using the PyTorch (Paszke et al., 2017) AdamW (Loshchilov & Hutter, 2019) optimizer with a learning rate of $41e-$ 4, a weight decay of $51e-$ 5 and a batch size of $1024$. We use the cross entropy loss with balanced class weights as the objective. The weights of the layer are normalized after each update and we use CLIP’s (Radford et al., 2021) temperature to scale the logits. We use the class-balanced accuracy on the validation set as the criterion for model selection and early stopping.
@@ 4.1 Datasets
Waterbirds (Wah et al., 2011b) is a common dataset for generalization and mitigating spurious correlations. It is created from CUB (Wah et al., 2011a), by grouping species of birds into two categories, landbirds and waterbirds, each being associated with a spurious correlation regarding its background, land, and water respectively.
@@
CelebA (Liu et al., 2015a) is a large-scale collection of celebrity images (over $200.000$), widely used in computer vision research. For generalization context, the setup (Liu et al., 2015b) consists of using the Blond_Hair attribute as the class label and the gender as the spurious feature.
@@
CivilComments (Borkan et al., 2019) is a large collection comprised of $1.8$ million online user comments. This dataset is used employed in NLP bias and fairness research concerning different social and ethnical groups.
@@
ImageNet (Deng et al., 2009) is a larger-scale popular dataset for image classification ($1000$ classes, with approx. $1300$ training samples and $50$ validation samples per class).
5 Evaluating spurious correlations
@@
For a proper quantitative evaluation of our proposed SCs, beyond the subjectivity of the qualitative aspects, we combine the concepts identified by WASP with different components: in Sec. 5.1 we generate samples exploiting the discovered concepts; in Sec. 5.2 we evaluate in scenarios lacking spurious correlation counterexamples; and in Sec. 5.3 we use a minimally SC-enhanced prompting for zero-shot prediction with a foundation model. A list of the extracted concepts for all datasets can be found in Appx. C.
@@
Table 3: Learning in the context of perfect spurious correlations. In the absence of samples that associate a class instance and concepts spuriously correlated with other classes, GroupDRO does not outperform the standard ERM. In contrast, our regularization based on the identified concepts consistently yields improvements (concerning worst group accuracy) over the considered baselines: ERM, GroupDRO, and the regularization with random causally unrelated concepts (obtained after the filtering in Step2a).
@@
Regularization w/ random concepts (class-neutral) 46.6$\pm$2.7 75.3$\pm$1.1 9.4$\pm$0.0 61.4$\pm$2.0 19.1$\pm$1.6 50.8$\pm$0.9
Regularization w/ SCs found by Lg 50.4$\pm$0.1 76.6$\pm$0.0 8.3$\pm$0.0 61.2$\pm$0.5 - -
Regularization w/ SCs found by WASP (ours) 57.9$\pm$0.3 79.8$\pm$0.1 10.4$\pm$0.5 62.0$\pm$1.8 31.3$\pm$0.7 57.5$\pm$0.4
5.1 ImageNet Spurious Correlations
We employ WASP to uncover a series of spurious correlations plaguing the decision-making process of OpenAI’s CLIP ViT-L/14 fine-tuned on ImageNet. As the SCs our method reveals lie beyond the restricted scope of the validation set, our sole recourse for illustrating and validating these findings at scale lies in generating synthetic images. To this end, we have employed a quantized version of FLUX.1-dev (Labs, 2024). To validate the impact of the SCs, we prompt the generative model to depict a chosen (correct) class alongside a non-ImageNet object that we identified as a positive SC for a different class.
The validation process is presented for three distinct scenarios in Tab. 1. Each scenario is defined by a correct class that is illustrated in the image, a concept (object, property, or activity) that is not causally tied to any class, and an absent class induced through the presence of the concept. We expect the classifier to predict this absent class, based on our scoring. We measure the impact of the concept by comparing the model’s ability to predict the correct class before and after its introduction. We generate $1000$ images for each scenario and we evaluate the model both in terms of accuracy and in terms of the frequency with which it predicts the induced class. Throughout all considered scenarios, we observe a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved, and a significantly increased likelihood of having the induced class predicted, even though it is not illustrated in any way, shape or form in the generated image.
Within the same context, we present a series of qualitative examples in Fig. 4. We emphasize that, even though throughout most of these samples, a single ImageNet-1k class is clearly depicted, the model chooses to ignore it and label the image as a completely different class, not illustrated at all in the image, solely based on the presence of a non-ImageNet object. We underline, by means of the results presented in Tab. 1, that the model is not fooled by artifacts in the generated images to predict randomly. We test the performance of the models on images featuring the correct class, without added objects. We observe this way that the model’s performance on the generated data is on par with the original performance of the model on these classes, validating that the generated images are not out of distribution. Furthermore, we show that the rate at which the induced class is predicted increases significantly, but more importantly, that the presence of the SC significantly increases the softmax activations of the model for the induced class (considering that it is spread across 1000 classes).
The model at hand is generally considered to be a robust state-of-the-art model, benefiting from ample pre-training, achieving an $82.4\%$ top-1 accuracy on the ImageNet validation set. We emphasize through our experiment that even under these reassuring circumstances, critical reasoning flaws can make their way through, in a production-ready model, undetected by validating and examining the model’s performance on held-out data. We further proceed to examine the impact of the SCs for which we have generated data on an exhaustive set of ImageNet-1k state-of-the-art models. We present the entire set of results in Appx. B, and the results for a selection of these models in Tab. 2.
Throughout this validation process we had to manually ensure that the data is generated correctly. Thus, we were forced to restrict ourselves to presenting only a few of our findings, ascribing a complete investigation of the impact of all SCs found to future work. We present a set of tables featuring the scores computed for different concepts with respect to different classes in Appx. C.
5.2 Training in a Fully Spurious Setup
We explore here an extreme setup, featuring no spurious correlation counterexample. In real world, this might be the case for most of the SCs, since usually this kind of correlations are generated by decisions in dataset acquisitions and filtering. Removing the minority groups from common robustness datasets, renders GroupDRO-like approaches completely ineffective, as their performance at best only matches the standard ERM (see Tab. 3). We use the employed SCs to impose a regularization upon the trained linear probes, learned on top of frozen embeddings from the FM. Intuitively, we constraint the weights to be equally distanced from the identified SCs. We formulate this as an MSE between the similarity of class weight $w_{k}$ with an SC $b$ and the average similarity of all class weights with $b$, these terms being scaled by $\tau$ (CLIP’s temperature):
@@ 5.2 Training in a Fully Spurious Setup
where $\mathcal{B}$ is the set of selected concepts and $\alpha=0.1$.
In Tab. 3, we present the results of linear probing with this loss, in the previously mentioned scenario, with no SC counterexamples, just 100% correlations between chosen tuples of concepts and classes. Through SC regularization, the learned classification weights are less reliant on the revealed SCs. The improvement in worst group accuracy shows the new representations are more robust and better aligned with the classification task, underling that WASP identifies concepts that are truly spuriously correlated with the classes. For a better comparison, we replicate the SC identification process of Lg (Zhao et al., 2024) in this scenario.
Table 4: Qualitative SCs examples, extracted on Waterbirds, CelebA and CivilComments datasets. See in red concepts that are off-topic, person names, or too related to the semantic content of the class, and in blue new concepts, that were not identified before. WASP, w.r.t. others, focus on learned SCs, discovering many new spuriously correlated concepts (and expressions, marked with …).
Waterbirds CelebA CivilComments
landbird waterbird blonde hair non-blonde hair offensive non-offensive
B2T forest, woods, tree, branch ocean, beach, surfer, boat, dock, water, lake - man, male *github - -
SpLiCE - - hairstyles, dolly, turban, actress, tennis, beard hairstyles, visor, amy, kate, fielder, cuff, rapper, cyclist - -
Lg forest, woods, rainforest, tree branch, tree beach, lake, water, seagull, pond …blonde hair, actress, model, woman long hair man, man wearing sunglasses, young man, black hair, actor - -
WASP (ours) forest…, bamboo…, ground, field, log, grass…, tree swimming…, water, lake, flying…, boat, lifeguard, pond - hat…, man…, actor, dark, person, large, shirt hypocrisy, troll, solly, hate allowing, work, made, talk
Table 5: Zero-shot prompting with Foundation Model (FM). To account the learned SCs, we explicitly introduce them in the prompting scheme, comparing with those discovered by competition. Using WASP spurious correlated concepts, the zero-shot performance (worst group accuracy) significantly improves across all datasets (image and text modalities).
@@
FM zero-shot 35.2 84.2 72.8 87.7 33.1 80.2
FM w B2T 48.1 86.1 72.8 88.0 - -
FM w SpLiCE - - 67.2 90.2 - -
FM w Lg 46.1 85.9 50.6 87.2 - -
FM w WASP 50.3 86.3 73.1 85.7 53.2 71.0
Table 6: Ablation. Following the zero-shot SC-augmented prompting setup, we variate the cut-off threshold for considering causally-unrelated concepts as spuriously correlated and also try to address the text-image modality gap.
@@
5.3 Spurious-aware zero-shot prompting
To further validate our identified spurious correlations, we follow Kim et al. (2024) and evaluate them in the context of a zero-shot classification task. We augment the initial, class-oriented prompt with the identified concepts through a minimal intervention (e.g. ’a photo of a {cls} in the {concept}’ (see Appx. A). For each class we create a prompt with each identified spurious correlation. When classifying an image, we take into account only the highest similarity among the prompts of a class (zero-shot with max-pooling over prompts). See in Tab. 5 how the SCs revealed by our method improve the worst group accuracy over the initial zero-shot baseline and other state-of-the-art solutions, in all the tested datasets. This highlights the relevance of the SCs automatically extracted by WASP.
5.4 Qualitative examples
We present in Tab. 4 the concepts identified as spuriously correlated with each class by WASP and competitors. Notice how our method discovers many new concepts (in blue) when compared with others. This is because our approach is fundamentally different, as it relies on the decision-making process of the model being investigated, diverging from current techniques oriented to validation set errors (B2T), or others that do data analysis over frozen concepts (SpLiCE, Lg). For CelebA-blonde hair, B2T and WASP do not found any SCs. This turn out to be an appropriate decision, since the presence of the feminine features do not incline the model towards one class or the other. See an exhaustive list of WASP revealed SCs (ImageNet-1k included) in Appx. C.
Ablations We validate several WASP decisions in Tab. 6, for zero-shot classification task, using prompts enhanced with SCs. The number of SCs per class turns out to be very important, taking too many adds noise to the prompts and lowers the performance. Nevertheless, dynamically choosing the threshold, as described in Sec. 3-Step2c., proves to be a good strategy for adapting the cut-off across classes. Following prior observations regarding the modalities gap between text and image embedding space (Liang et al., 2022; Zhang et al., 2023a), we subtract half of the gap from the embeddings and re-normalize them, ending in marginally lower performance w.r.t. not addressing the gap.
@@ 6 Conclusions
We introduce WASP, a method that automatically identifies SCs through a weight-space and filtration approach, that can be scaled to large datasets. By leveraging foundation models’ embedding space topologies, we identify concepts that significantly influence class predictions, while filtering undesirable concepts. We apply this method to multiple modalities on datasets like Waterbirds, CelebA, CivilComments, and ImageNet. We validate that the proposed SCs are useful by employing them in different methods for improved generalization. This includes a) a novel SC-based regularization that works even under $100\%$ correlation of concepts and classes, where previous methods fail b) using better SC-aware zero-shot promoting. Using WASP we identify SCs in ImageNet and generate synthetic images that are able to fool strong pre-trained methods, validating the relevance of the identified SCs in larger datasets. This method opens the door for automatically finding SCs in large settings without explicit human intervention.